# Towards Robust Recommendation: A Review and an Adversarial Robustness Evaluation Library

### v1

@@ 4.2.3 Trust-aware recommender
For example, Zhang et al. (2017)[55] calculated user trustworthiness based on social data and used trustworthiness values to weight-aggregate neighbor ratings, thereby improving recommendation accuracy and robustness. Quan et al. (2023) designed the Preference-Guided Denoising Method with Social Relationships (GDMSR)[56]. GDMSR models confidence in social relationships and provides a denoised and richer social graph to enhance user preference learning. To mitigate noisy social relationships, GDMSR also incorporates a self-correcting curriculum learning module and an adaptive denoising strategy, both of which contribute to improved trust in high-degree samples. Wang et al. (2023) introduced the Denoised Self-Augmented Learning (DSL) paradigm[57]. DSL starts with self-supervised denoising, preserving useful social relationships. It then employs a dual-view graph neural network to encode latent representations from user social and interaction graphs. A learnable similarity mapping function aligns these two views, achieving robust recommendations.
@@
For example, Jia et al. (2013)[58] assessed the credibility of user ratings based on the reliability of item recommendations, score similarity, and user trustworthiness. They combined a trust computation model with traditional CF methods, selecting a reliable set of neighbors to generate recommendations for the target user. Dongyan and Fuzhi (2014)[59] employed entropy theory and the concept of density-based local outlier factors to calculate user suspicion levels. They introduced a multidimensional trust model that measured user trust attributes from various perspectives. This model, which includes user suspicion levels, was then combined with baseline estimation methods to create the Trusted Neighborhood Model. Gao et al. (2014)[60] proposed a user relationship-based method to enhance the robustness of recommender systems. They first constructed user models using three common user relationships (interest similarity, score similarity, and trust relationships) and used density clustering algorithms to detect spam users. Then, they introduced a detection-based approach to calculate item similarity and applied a typical item-based k-Nearest Neighbors (kNN) CF method for rating prediction. Experimental results demonstrated the effectiveness of this approach in detecting spam users and improving recommender system robustness. Yi and Zhang (2016)[61] introduced the RRM-SUMMT method. They manually labeled suspicious users and trained a relevance vector machine classifier to determine user suspicion levels. Subsequently, they constructed a multidimensional trust model using user rating data and combined MF with neighborhood models for recommendation.
4.3 Hybrid method and others
To simultaneously enhance the accuracy and robustness of recommender systems, some studies have combined detection and robustness algorithms for recommendation.
@@ 4.3 Hybrid method and others
5 Non-adversarial robustness
5.1 Definition And Taxonomy
@@ 5.2 Data sparsity
For instance, Li and She (2017) proposed a Bayesian generative model based on Collaborative Variational Autoencoders (CVAE)[3] that integrates ratings and recommendation content in multimedia scenarios. It learns deep latent representations from content data in an unsupervised manner and captures implicit relationships between items and users from both content and ratings. Experiments showed that CVAE alleviates data sparsity and cold-start problems. Yang et al. (2022)[4] combined knowledge graph data with contrastive learning to leverage the rich data in knowledge graphs to learn the true preferences of users with sparse interactions, thereby improving recommendation accuracy. Chen et al. (2019)[5] improved recommendation performance in hash-based recommender systems by transforming Locality-Sensitive Hashing (LSH) into Multiple Probing LSH and integrating it into the auxiliary recommendation process, using collaborative recommendation principles to enhance performance in cases of sparse interactions. Similarly, Chen et al. (2019)[65] proposed a method that combines LSH with Hybrid CF (HCF) to enhance recommendation performance in scenarios with sparse interactions.
5.3 Natural Noise in Implicit Feedback
@@
Building upon this, the denoising of implicit feedback recommender systems holds significant importance. According to our research findings, we classify the denoising task of implicit feedback into two categories: sample selection methods and sample re-weighting methods.
@@
Dataset #Users #Itmes #Ratings Rating range Sparsity Average action per user Average action per item Timestamp Social data
@@
Dataset #Users #Itmes #Interactions Sparsity Average action per user Average action per item Timestamp Social data
@@
5.3.1 Sample selection methods
For instance, Gantner et al. (2011) introduced the Bayesian Personalized Ranking framework WBPR[66], which is applicable in scenarios with uneven sampling. The core idea behind WBPR is the assumption that popular but non-interacted items are likely to be genuine negative items. Zhang et al. (2013)[67] proposed a hypothesis that during training, the ranking of any unobserved item should not be higher than that of any observed positive training item. Based on this hypothesis, the authors introduced DNS, a method that dynamically selects negative training samples from the ranking lists generated by the current prediction model. It employs an n-degree polynomial weighted function to select n+1 samples and uses them to calculate the probability of each item being sampled. Wang et al. (2021)[68] presented a novel iterative re-ranking framework, IRBPR, designed to alleviate noise in implicit feedback. The iterative re-ranking module employs self-training principles to dynamically generate pseudo-labels based on user preferences. These refreshed labels are then used to train downstream modules for the recommendation task. Experimental results demonstrate that IRBPR effectively mitigates noise issues. Zhu et al. (2022)[69] pointed out that previous work lacked sufficient robustness to handle false negative samples during training. Therefore, the authors proposed GDNS, which designs an expected gain sampler. This sampler dynamically guides negative sample selection based on the expected preference gap between positive and negative samples during user training. This gain-adjusted negative sampler can effectively identify false negative samples and further reduce the risk of introducing false negative instances.
5.3.2 Sample re-weighting methods
While sample selection methods can address the noise issue to some extent in implicit feedback recommender systems, they may unavoidably remove genuine samples due to the pursuit of high recall rates. Additionally, the performance of sample selection methods heavily relies on the data distribution. To address the aforementioned challenges, a series of denoising methods based on sample re-weighting have been proposed. Sample re-weighting methods distinguish between clean and noisy samples based on sample loss values, assigning lower weights to noisy samples during model training[70].
5.4 Data imbalance
6 Dataset for robust recommendation
We investigated and compiled commonly used datasets for evaluating the robustness of recommendation systems. These datasets were categorized into explicit feedback datasets and implicit feedback datasets based on whether they contain explicit user preference information. The statistical information can be found in Table II and Table III.
6.1 Explicit feedback dataset
@@
The Book Crossing Dataset [80]is an online book review data set collected and released by a group of individuals. The Book-Crossing dataset comprises book ratings, book titles, and user ages.
6.2 Implicit feedback dataset
@@
Figure 6: The overall framework of ShillingREC
7 Evaluation Metrics
Based on varying task requirements, we categorize recommender system evaluation metrics into four classes: rating prediction metrics, top-k recommendation metrics, and robustness metrics. In this subsection, we will provide a detailed overview of each category of metrics.
7.1 Rating prediction metrics
@@ 7.2 Top-k recommendation metrics
$\mathrm{HR}@N=\frac{\mathrm{hits}}{\left|\mathcal{U}\right|}$ (9)
@@ 7.3 Robustness metrics
, where $\hat{r}{ui}$ represents the predicted score by the recommendation model for the user-item pair before the attack, and $\hat{r}{ui}^{\prime}$ represents the predicted score by the recommendation model for the user-item pair after the attack.
@@
TABLE IV: Existing recommender system libraries.
@@
Library Language Framework Application Description Release time #Star999#Star represents the number of stars for the project on the GitHub website, with data collection ending on April 21, 2024.
@@
LibRec [^1] https://github.com/guoguibing/librec Java DL4J Academic An open-source Java-based library designed to efficiently address rating and ranking challenges within recommender systems. This library facilitates rapid resolution and supports deployment across multiple platforms, including compatibility with Hadoop and Spark frameworks. 2014 3.2k
@@
Recommenders [^2] https://github.com/recommenders-team/recommenders Python Tensorflow Academic This repository encompasses illustrative examples and optimal methodologies for constructing recommender systems, presented in the form of Jupyter notebooks. 2018 18k
@@
EasyRec [^3] https://github.com/alibaba/EasyRec Python tensorflow Industry It supports training across multiple platforms and data sources, allowing the exploration of arbitrary parameters for enhanced flexibility in model optimization. 2020 1.5k
@@
ReChorus [^4] https://github.com/THUwangcy/ReChorus/tree/master Python pytorch Academic A rapid evaluation framework tailored for top-k recommendation models, featuring concise code and bolstered by multi-threaded batch preparation capabilities. 2020 490
@@
RecBole [^5] https://github.com/RUCAIBox/RecBole Python pytorch Academic This framework supports various types of recommendation tasks, including General Recommendation, Sequential Recommendation, Context-aware Recommendation, and Knowledge-based Recommendation. It accommodates over 40 datasets and includes preprocessing tools for efficient data handling. 2020 3.2k
@@
RecAd [^6] https://github.com/gusye1234/recad Python pytorch Academic An open benchmark for evaluating and testing recommender attack and defense strategies, supporting the assessment of common attack models on recommender systems and the evaluation of recommender models’ robustness. 2023 21
@@
ARLib [^7] https://github.com/CoderWZW/ARLib?tab=readme-ov-file Python pytorch Academic The primary focus lies in addressing attacks and defenses in top-k recommendation tasks. This involves adapting various adversarial attack methods to implicit feedback data and employing pairwise strategies to optimize recommendation models. 2023 74
@@
ShillingREC [^8] https://github.com/chengleileilei/ShillingREC Python pytorch Academic The framework supports adversarial attacks on both implicit and explicit feedback data, enabling the evaluation of different categories of recommendation models based on the data type and the nature of the recommendation task. 2024
In order to address the aforementioned issues concerning various types of recommender system attacks and existing defense strategies against such attacks, we aim to establish a unified evaluation framework to systematically assess them. This approach facilitates the analysis of the characteristics and applicability of different methods and provides a fair competition platform for various attack and defense techniques. To tackle this challenge, we conducted an investigation on open-source recommender system frameworks, as shown in Table IV. This investigation includes evaluation frameworks tailored for recommendation tasks, as well as those designed specifically to assess recommender system attacks. While several evaluation frameworks for attack, such as RecAd [90] and ARLib [91], already exist, RecAd primarily targets attacks on rating data, lacking adaptation for implicit feedback data. In contrast, ARLib is designed for implicit feedback data, converting some rating data-based attack methods to attacks on implicit feedback data and evaluating them in ranking tasks. However, this setting may not be equitable for certain attack methods designed for rating data.
8 Adversarial robustness evaluation Library and Experiment
8.1 The Library: ShillingREC
@@
To address the aforementioned issues, we propose an adversarial robustness evaluation library named ShillingREC, which supports various attack modes and recommendation tasks. The overall framework of ShillingREC is illustrated in Figure 6. ShillingREC can flexibly design attack modes based on different data types and recommendation task types. Specifically, as shown in Table I, we categorize shilling attacks into attacks on rating data and attacks on implicit feedback data. For attacks on rating data, we evaluate them on rating prediction tasks and ranking tasks based on pointwise optimization strategies. For attacks on implicit feedback data, we evaluate them on ranking tasks based on pairwise optimization strategies. Moreover, ShillingREC supports multiple commonly used datasets and is easily extensible. Integrates well-designed execution processes and comprehensive persistence strategies for data preprocessing, poisoning attacks, and recommendation model training.
@@
Our library consists of five main components: the Configuration Module, the Data Module, the Model Module, the Execution Module, and the Logger Module. Below, we provide a brief overview of the designs of these five modules.
8.1.1 Configuration Module
@@ 8.1.2 Data Module
8.1.3 Model Module

### v3

@@ IV-B3 Trust-aware Recommender
For example, Quan et al. (2023) designed the Preference-Guided Denoising Method with Social Relationships (GDMSR)[58]. GDMSR models confidence in social relationships and provides a denoised and richer social graph to enhance user preference learning. To mitigate noisy social relationships, GDMSR also incorporates a self-correcting curriculum learning module and an adaptive denoising strategy, both of which contribute to improved trust in high-degree samples. Wang et al. (2023) introduced the Denoised Self-Augmented Learning (DSL) paradigm[59]. DSL starts with self-supervised denoising, preserving useful social relationships. It then employs a dual-view graph neural network to encode latent representations from user social and interaction graphs. A learnable similarity mapping function aligns these two views, achieving robust recommendations.
@@
For example, Jia et al. (2013) [60] assessed the credibility of user ratings based on item recommendation reliability, score similarity, and user trustworthiness. They combined a trust computation model with traditional CF methods, selecting a reliable set of neighbors to generate recommendations. Dongyan and Fuzhi (2014) [61] used entropy theory and density-based local outlier factors to calculate user suspicion levels. They proposed a multidimensional trust model to measure user trust from multiple perspectives, which was then combined with baseline estimation methods to form the Trusted Neighborhood Model. Gao et al. (2014) [62] proposed a user relationship-based method to enhance robustness. They constructed user models based on three types of relationships (interest similarity, score similarity, and trust), used density clustering to detect spam users, and introduced a detection-based item similarity measure within an item-based kNN CF framework. Experiments demonstrated its effectiveness in spam detection and robustness improvement. Yi and Zhang (2016) [63] introduced RRM-SUMMT, manually labeled suspicious users, and trained a relevance vector machine to determine suspicion levels. They then built a multidimensional trust model from rating data and combined MF with neighborhood models for recommendation.
IV-C Hybrid Method and Others
To simultaneously enhance the accuracy and robustness of recommender systems, some studies have combined detection and robust algorithms for recommendation.
@@ IV-C Hybrid Method and Others
V Non-adversarial Robustness
V-A Definition and Taxonomy
@@ V-B Data Sparsity
For example, Li and She (2017) proposed a Bayesian generative model based on Collaborative Variational Autoencoders (CVAE)[2] that integrates ratings and recommendation content in multimedia scenarios. It learns deep latent representations from content data in an unsupervised manner and captures implicit relationships between items and users of both content and ratings. Experiments showed that CVAE alleviates data sparsity and cold start problems. Chen et al. (2019)[4] improved the performance of recommendations in hash-based recommender systems by transforming locality sensitive hammering (LSH) into multiple probing LSH and integrating it into the auxiliary recommendation process, using collaborative recommendation principles to improve performance in cases of sparse interactions. Yang et al. (2021) propose an enhanced graph learning network for CF called EGLN [67]. EGLN iterates between enhanced graph learning and node embedding learning, adding potential links absent in the input graph. By maximizing mutual information between local and global representations, it strengthens user-item relationships and improves performance on sparse data. Yang et al. (2022)[3] combined knowledge graph data with contrastive learning to take advantage of the rich data in knowledge graphs to learn the true preferences of users with sparse interactions, thus improving the precision of the recommendation. Similarly, Du et al. (2023)[68] proposed a metalearning framework based on knowledge graphs, MetaKG. Using a collaborative-aware meta-learner to capture user preferences and a knowledge-aware meta-learner to capture entity knowledge, MetaKG integrates an adaptive task scheduler to filter information tasks, ensuring recommendation performance in scenarios of data sparsity and cold starts.
V-C Natural Noise in Implicit Feedback
@@
Building upon this, the denoising of implicit feedback recommender systems holds significant importance. According to our research findings, we classify the denoising task of implicit feedback into two categories: sample selection methods and sample reweighting methods.
V-C1 Sample Selection Methods
For instance, Gantner et al. (2011) introduced the Bayesian Personalized Ranking framework WBPR[69], applicable to scenarios with uneven sampling. The core idea of WBPR is that popular but non-interacted items are likely genuine negatives. Zhang et al. (2013) [70] hypothesized that during training, any unobserved item should not rank higher than any observed positive item. Based on this, they proposed DNS, which dynamically selects negative samples from ranking lists generated by the current model. It uses an n-degree polynomial weighted function to select n+1 samples and compute sampling probabilities. Wang et al. (2021) [71] presented an iterative re-ranking framework, IRBPR, to alleviate noise in implicit feedback. The re-ranking module applies self-training to dynamically generate pseudo-labels based on user preferences, which are then used to train downstream recommender modules. Experiments show IRBPR effectively mitigates noise. Zhu et al. (2022) [72] noted that previous work lacked sufficient robustness to handle false negatives during training. To address this, they proposed GDNS, which employs an expected gain sampler to dynamically guide negative sample selection based on the expected preference gap between positive and negative items. This gain-adjusted sampler effectively identifies false negatives and reduces the risk of introducing them.
V-C2 Sample Reweighting Methods
While sample selection methods can address the noise issue to some extent in implicit feedback recommender systems, they may unavoidably remove genuine samples due to the pursuit of high recall rates. Additionally, the performance of sample selection methods heavily relies on the data distribution. To address the aforementioned challenges, a series of denoising methods based on sample reweighting have been proposed. Sample reweighting methods distinguish between clean and noisy samples based on sample loss values, assigning lower weights to noisy samples during model training[73].
V-D Data Imbalance
@@
Dataset #Users #Itmes #Ratings Rating range Sparsity Avg Actions/User Avg Actions/Item Timestamp Social data
@@
Dataset #Users #Itmes #Interactions Sparsity Avg Actions/User Avg Actions/Item Timestamp Social data
@@
VI Dataset for Robust Recommendation
We investigated and compiled commonly used datasets for evaluating the robustness of recommender systems. These datasets were categorized into explicit feedback datasets and implicit feedback datasets based on whether they contain explicit user preference information. The statistical information can be found in Table II and Table III.
VI-A Explicit Feedback Dataset
@@
The Book-Crossing Dataset [83] is an online book review dataset collected and released by a group of individuals, comprising book ratings, titles, and user ages.
VI-B Implicit Feedback Dataset
@@
VII Evaluation Metrics
Based on varying task requirements, we categorize recommender system evaluation metrics into three classes: rating prediction metrics, top-k recommendation metrics, and robustness metrics. In this subsection, we will provide a detailed overview of each category of metrics.
VII-A Rating Prediction Metrics
@@ VII-B Top-k Recommendation Metrics
$\mathrm{HR}@K=\frac{\mathrm{hits}}{\left|\mathcal{U}\right|}$ (9)
@@ VII-C Robustness Metrics
, where $\hat{r}_{ui}$ represents the predicted score by the recommendation model for the user-item pair before the attack, and $\hat{r}_{ui}^{\prime}$ represents the predicted score by the recommendation model for the user-item pair after the attack.
The T-metric evaluates ranking-based performance specifically on the target items of the attack. A lower T-metric value indicates a lower attack success rate and thus stronger robustness of the model against shilling attacks. Common examples include T-NDCG and T-HR.
@@
TABLE IV: Existing Recommender System Libraries
@@
Library Language Framework Description Release time #Star999#Star represents the number of stars for the project on the GitHub website, with data collection ending on April 21, 2024.
@@
LibRec [^1] https://github.com/guoguibing/librec Java DL4J An open-source Java library designed for rating and ranking tasks in recommender systems, supporting rapid deployment on platforms like Hadoop and Spark. 2014 3.2k
@@
Recommenders [^2] https://github.com/recommenders-team/recommenders Python Tensorflow This repository encompasses illustrative examples and optimal methodologies for constructing recommender systems, presented in the form of Jupyter notebooks. 2018 18k
@@
EasyRec [^3] https://github.com/alibaba/EasyRec Python Tensorflow It supports training across multiple platforms and data sources, allowing the exploration of arbitrary parameters for enhanced flexibility in model optimization. 2020 1.5k
@@
ReChorus [^4] https://github.com/THUwangcy/ReChorus/tree/master Python Pytorch A rapid evaluation framework tailored for top-k recommendation models, featuring concise code and bolstered by multi-threaded batch preparation capabilities. 2020 490
@@
RecBole [^5] https://github.com/RUCAIBox/RecBole Python Pytorch It supports general, sequential, context-aware, and knowledge-based recommendation tasks, along with 40+ datasets and preprocessing tools. 2020 3.2k
@@
RecAd [^6] https://github.com/gusye1234/recad Python Pytorch An open benchmark for evaluating recommender attacks and defenses, supporting robustness assessment of models against common attacks. 2023 21
@@
ARLib [^7] https://github.com/CoderWZW/ARLib?tab=readme-ov-file Python Pytorch Focuses on attacks and defenses in top-k recommendation by adapting adversarial attack methods to implicit feedback. 2023 74
@@
ShillingREC [^8] https://github.com/chengleileilei/ShillingREC Python Pytorch The framework supports adversarial attacks on both implicit and explicit feedback data, enabling the evaluation of different categories of recommendation models based on the data type and the nature of the recommendation task. 2024
VIII Adversarial Robustness Evaluation Library and Experiment
In this chapter, we aim to fairly evaluate poisoning attack models and recommendation models under adversarial robustness scenarios to analyze the performance and applicability of different models. Therefore, we investigated open-source recommendation system frameworks, as shown in Table  IV. Some open-source frameworks support the evaluation of poisoning attack and defense models, such as RecAd [93] and ARLib [94]. However, these frameworks are not without limitations. RecAd only supports poisoning attacks on rating datasets, lacking adaptation for implicit data. In contrast, ARLib is designed for implicit data but forces poisoning attack methods for rating data to be adapted for implicit feedback data during evaluation. This setup is unfair to the former, as rating-based attack methods require generating specific fake ratings, while implicit data attack methods only need to generate fake interaction data.
Figure 6: The overall framework of ShillingREC.
VIII-A The Library: ShillingREC
@@
To address the aforementioned issues, we propose ShillingREC, an adversarial robustness evaluation library that supports multiple attack modes and recommendation tasks. The overall framework of ShillingREC is illustrated in Figure 6. ShillingREC enables poisoning attacks on both rating and implicit data formats. For rating data, it injects fake rating data and evaluates model performance using rating prediction tasks. For implicit feedback data, it injects fake interaction data and evaluates performance through top-k recommendation tasks. This design allows for a fair comparison of different poisoning attack models within their respective areas of strength. Additionally, ShillingREC supports various commonly used datasets, offering excellent scalability. It also provides efficient execution pipelines and comprehensive persistence strategies for data processing, poisoning attacks, and recommendation model training.
@@
ShillingREC consists of five main components: Configuration Module, Data Module, Model Module, Execution Module, and Logger Module.
VIII-A1 Configuration Module
@@ VIII-A2 Data Module
TABLE V: Evaluation Results of Explicit Feedback Data
Attack Model Rec Model ML-100K ML-1M CiaoDVD Epinions
RMSE MAE PS RMSE MAE PS RMSE MAE PS RMSE MAE PS
None MF 1.1778 0.9823 0 1.1113 0.9276 0 1.5482 1.3238 0 1.3183 1.1490 0
NeuMF 0.9556 0.7520 0 0.8996 0.7071 0 0.9602 0.7290 0 1.0445 0.7840 0
LightGCN 0.9773 0.7678 0 0.8785 0.6894 0 1.2865 0.9809 0 1.1404 0.8822 0
SGL 0.9974 0.7820 0 0.8955 0.7015 0 1.3868 1.0554 0 1.2133 0.9405 0
Random Attack MF 1.1693 0.9779 1.6996 1.1090 0.9269 1.2985 1.4541 1.2490 2.2385 1.3213 1.1569 0.6326
NeuMF 0.9577 0.7548 1.9365 0.8975 0.7069 1.7194 0.9553 0.7270 0.8949 1.0570 0.7939 1.1908
LightGCN 0.9529 0.7519 1.8517 0.8723 0.6855 1.3987 1.2108 0.9346 1.4879 1.1036 0.8617 2.2824
SGL 0.9667 0.7618 2.0922 1.0701 0.8482 2.4756 1.3301 1.0234 1.4936 1.1678 0.9130 2.6397
Average Attack MF 1.1831 0.9864 1.8206 1.1129 0.9294 1.4259 1.4535 1.2483 2.2080 1.3172 1.1489 0.4278
NeuMF 0.9619 0.7567 2.2095 0.9009 0.7037 2.2377 0.9545 0.7245 0.8340 1.0558 0.7864 1.0411
LightGCN 0.9595 0.7568 3.0650 0.8729 0.6860 2.4352 1.2141 0.9365 1.4376 1.1006 0.8555 2.3379
SGL 0.9745 0.7674 3.3349 0.8814 0.6926 2.9289 1.0204 1.3297 1.5289 1.1627 0.9058 2.5788
LoveHate Attack MF 1.1790 0.9829 0.5294 1.1111 0.9271 0.8449 1.5052 1.2866 0.5978 1.3198 1.1503 0.5983
NeuMF 0.9575 0.7495 4.1547 0.8895 0.6968 3.9182 0.9578 0.7292 1.8924 1.0496 0.7861 3.6649
LightGCN 0.9784 0.7689 3.3856 0.8750 0.6857 5.2998 1.3038 0.9811 2.4647 1.1202 0.8647 4.5518
SGL 0.9886 0.7743 5.1468 0.8885 0.6964 4.2893 1.4142 1.0684 2.3064 1.1945 0.9249 3.8280
AUSH MF 1.1845 0.9875 1.7807 1.1113 0.9278 1.7045 1.4983 1.2795 2.2927 1.3163 1.1478 0.4278
NeuMF 0.9631 0.7577 2.3155 0.8912 0.6967 2.1448 0.9679 0.7335 0.8692 1.0597 0.7915 0.8401
LightGCN 0.9673 0.7586 5.8488 0.8808 0.6906 2.3850 1.2566 0.9558 1.9112 1.1722 0.8992 3.4358
SGL 0.9999 0.7835 3.9206 0.8979 0.7030 3.3751 1.3879 1.0542 2.3133 1.2271 0.9439 4.2808
TABLE VI: Evaluation Results of Implicit Feedback Data(K=50)
Attack Model Rec Model Gowalla LastFM ML-1M CiaoDVD
NDCG HR T-NDCG T-HR NDCG HR T-NDCG T-HR NDCG HR T-NDCG T-HR NDCG HR T-NDCG T-HR
None MF 0.0303 0.1491 0 0.2000 0.0461 0.4133 0.0002 0.0014 0.1346 0.8027 0 0 0.0359 0.2900 0.0114 0.0433
NeuMF 0.0668 0.3174 0.0003 0.0038 0.0676 0.5249 0 0 0.1575 0.8800 0 0 0.0938 0.5628 0 0
LightGCN 0.0821 0.3621 0.0001 0.0016 0.0674 0.5130 0 0 0.1612 0.8862 0 0 0.0874 0.5195 0.0044 0.0130
CDAE 0.0226 0.1553 0 0 0.0614 0.5151 0 0 0.1618 0.8966 0 0 0.0665 0.4199 0 0
SGL 0.0799 0.3599 0 0.0002 0.0557 0.4765 0.0005 0.0021 0.1609 0.8776 0 0 0.0999 0.5628 0.0192 0.0563
SGDL 0.0642 0.0002 0 0 0.0674 0.5207 0 0 0.1584 0.8529 0 0 0.0831 0.5043 0.0011 0.0043
Random Attack MF 0.0278 0.1397 0.0146 0.0861 0.0384 0.3726 0.0027 0.0105 0.1368 0.8000 0 0 0.0415 0.3074 0.0196 0.0649
NeuMF 0.0616 0.2977 0.3026 0.5843 0.0577 0.5081 0.4606 1 0.1549 0.8662 0.0003 0.0018 0.0826 0.5498 0.9089 0.9307
LightGCN 0.0796 0.3556 0.0173 0.0780 0.0663 0.5109 0 0 0.1593 0.8764 0 0 0.0975 0.5649 0.0369 0.1147
CDAE 0.0184 0.1457 0.9977 1 0.0554 0.5039 0.7955 0.8561 0.1622 0.8995 0.0001 0.0017 0.0637 0.4091 0.2332 0.9978
SGL 0.0772 0.3462 0.0291 0.1663 0.0605 0.4940 0.0035 0.0119 0.1608 0.8751 0.0001 0.0003 0.0931 0.5693 0.0028 0.0130
SGDL 0.0704 0.3433 0.0701 0.2398 0.0668 0.5158 0 0 0.1587 0.8596 0 0.0005 0.0831 0.5152 0.0024 0.0087
Bandwagon Attack MF 0.0278 0.1368 0.0385 0.1663 0.0361 0.3565 0.0058 0.0309 0.1339 0.7961 0.0001 0.0005 0.0403 0.3030 0.0169 0.0628
NeuMF 0.0624 0.2984 0.1650 0.4381 0.0615 0.5172 0.7368 1 0.1560 0.8672 0.0002 0.0017 0.0834 0.5519 0.6369 0.9199
LightGCN 0.0791 0.3543 0.0621 0.2202 0.0670 0.5235 0 0 0.1601 0.8813 0 0 0.0987 0.5801 0.0295 0.0779
CDAE 0.0190 0.1462 0.7748 1 0.0593 0.5193 0.0024 0.0189 0.1607 0.8963 0.0001 0.0008 0.0645 0.4134 0.2495 0.9978
SGL 0.0780 0.3534 0.0162 0.0700 0.0556 0.4814 0.0022 0.0077 0.1612 0.8754 0.0005 0.0039 0.0962 0.5563 0.0008 0.0043
SGDL 0.0693 0.3424 0.0761 0.1851 0.0675 0.5200 0 0 0.1591 0.8576 0 0.0005 0.0837 0.5130 0.0022 0.0087
AUSH MF 0.0391 0.1969 0.0646 0.1558 0.0461 0.4456 0.5428 0.7733 0.1324 0.8022 0.1348 0.4569 0.0599 0.3939 0.0349 0.0801
NeuMF 0.0641 0.3100 0.1674 0.3684 0.0563 0.5165 0.9984 1 0.1579 0.8756 0.0193 0.0632 0.0824 0.5844 0.7730 0.8247
LightGCN 0.0802 0.3596 0.1634 0.3040 0.0565 0.5130 0.9977 0.9993 0.1616 0.8872 0.0006 0.0017 0.0874 0.5260 0.0077 0.0260
CDAE 0.0178 0.1426 0.9998 1 0.0544 0.5144 0.9296 0.9544 0.1599 0.8985 0.0005 0.0025 0.0636 0.4048 0.2016 0.9957
SGL 0.0797 0.3578 0.0494 0.1207 0.0532 0.4709 0.6427 0.8182 0.1610 0.8734 0.0004 0.0013 0.0963 0.5844 0.0041 0.0195
SGDL 0.0693 0.3487 0.1420 0.2776 0.0556 0.5039 0.9984 1 0.1595 0.8543 0.0033 0.0183 0.0820 0.5087 0.0077 0.0325
RevAdv MF 0.0349 0.1676 0.0393 0.1632 0.0436 0.4126 0.2504 0.6912 0.1361 0.7926 0 0.0002 0.0180 0.1883 0.0005 0.0022
NeuMF 0.0654 0.3100 0.1672 0.5387 0.0616 0.5123 0.4192 1 0.1573 0.8754 0.0014 0.0097 0.0967 0.5758 0.2764 0.7208
LightGCN 0.0820 0.3637 0.0491 0.1690 0.0615 0.5102 0.7367 1 0.1607 0.8897 0.0021 0.0171 0.0851 0.5216 0.1734 0.6883
CDAE 0.0729 0.3364 0.2083 0.6596 0.0595 0.5060 0.1476 0.7895 0.1596 0.8983 0.0105 0.1032 0.0649 0.4156 0.1952 0.9935
SGL 0.0796 0.3612 0.0150 0.0773 0.0572 0.4772 0.1835 0.7460 0.1609 0.8798 0.0024 0.0173 0.0996 0.5801 0.0847 0.3506
SGDL 0.0699 0.3438 0.0279 0.1477 0.0635 0.4989 0.6017 1 0.1596 0.8571 0 0 0.0829 0.5130 0.1742 0.7100
VIII-A3 Model Module